### v1

@@ 4 Experimental Results
The ideal public bidding network consists of nodes uniformly connected and characterized by the nonexistence of strong and connected communities, leading to a challenging trade-off: extracting connections from this set and retaining those that represent a high-level interaction between companies without overemphasizing what could be simple statistical deviations. In this scenario, Heron’s Information Coefficient is useful due to its quantification of structural imbalance in dynamic networks, revealing signs of the coordinated behavior typical of collusion. Here we use the disparity filter algorithm [26] alongside the maximization of HIC to highlight the central players in the BID network.
The idea of the disparity filter is to associate to each network’s link a weight $\alpha_{ij}$ that represents a significant level provided by the distribution of the local uniform randomness null hypothesis [26]. Relevant links are chosen so that $\alpha_{ij}<\alpha$, $\alpha\in[0,1]$ being a fixed significance level. Instead of choosing the significance levels arbitrarily, we maximize HIC when $\alpha$ varies. Disconnected companies are excluded from the resulting network. The process is repeated until reaching a significance level as small as desired. The selection of the most relevant group is made iteratively, starting with the original network with a significance level equal to 1. A new subgraph is obtained at each step by removing edges possessing a higher significant level. The goal is to find a lower significance level for each step, resulting in a higher HIC. By definition, this happens when the topologies of the active, inactive and original networks are close to equidistant.
Each company that participated in the bidding process is a node connected by links weighted by the number of times they have competed in the same bidding process. The way the nodes interact and connect provides essential information about the system. In particular, information about their connectivity patterns represents a valuable tool on how different dynamics occur in the system. For example, the speed at which information travels within a cohesive community is much greater than in a more sparse one. As such, let $G$ be the original weighted network after applying the disparity filter, for a fixed ${\alpha}\in[0,1]$, and $G_{a}^{\alpha}$ be the network obtained from $G$ by removing links whose weight is greater than or equal to $\alpha$. As such, $G_{i}^{\alpha}=G-G_{a}^{\alpha}$ is the complement of $G_{a}^{\alpha}$ to $G$. We define $\alpha_{T}$ as the smallest value of $\alpha$ such that ${\cal H}_{e}(G,G_{a}^{\alpha},G_{i}^{\alpha})$ is as large as possible. This optimal threshold $\alpha_{T}$ thus identifies the most informative backbone of the network, balancing cohesion and separation between active and inactive subgraphs. In doing so, it reveals the most representative organization of competitive interactions among companies.
All the results on this section are based on the extraction of the network’s backbone with this method. The source code for the characterization of our coefficient’s behavior and other experiments can be found on GitHub.
@@ 4.1 Artificial bidding network
We tested our method on artificial procurement networks generated using the Bernoulli Weighted Random Network (BWRN) model [27], which combines the Stochastic Block Model (SBM) for group formation with hierarchical structures for management layers. This generator creates synthetic networks that are statistically equivalent substitutes for real networks, allowing for research on covert networks. Its validity has been established through comparison with the publicly available Sicilian Mafia network [24], which is a covert network spanning phone calls and in-person meetings extracted from court proceedings. These synthetic networks thus replicate behaviors characteristic of real colluding agents while addressing a critical challenge in covert network research: data availability [28]. We enhanced realism by incorporating successful bids to better simulate procurement networks with known corrupt agents. Then, we iterated the HIC filtering process until network sparsity prevented further analysis.
@@
To validate robustness, we tested the method across varying parameters and 100 random seeds, iterating until all HIC values reached zero, that is, the network dissolved. The results were consistent, as seen in 7. With a lower percentage of corrupt agents, the method stopped faster. Networks with lower corruption levels (5%, 10%) maintained higher accuracy throughout, generally staying above 0.8. Furthermore, the method showed great final accuracy for networks containing up to 30% corrupt agents, with the final iteration successfully identifying at least one corrupt agent in most cases. Performance remains stable across network scales, though optimization would be necessary for graphs exceeding 10,000 nodes.
@@
In summary, our method is computationally cheap, converges fast and shows promising results for detecting corrupt agents in bidding processes. The Heron Information Coefficient effectively identifies corrupted agents in early iterations, particularly at lower corruption levels. While accuracy becomes less stable at corruption levels exceeding 50% - an expected limitation since anomaly detection inherently struggles when anomalous nodes constitute the majority - the method remains valuable. Even after six iterations in high-corruption scenarios, it substantially reduces network size while retaining corrupt nodes. As such, our results suggest that the Heron coefficient can serve as a practical aid for corruption inspection, even in resource-constrained environments with limited detection capabilities.
@@ 4.2 The BID Network
Brazil is a country of continental dimensions, the fifth largest in the world and the sixth in terms of population [29] with considerable regional and social inequalities [10, 9, 4]. As determined by the country’s Constitution, the provision of services in healthcare and education is a citizen’s right, met with resources collected through taxation at the federal, state and municipal levels [4, 6, 10, 7]. Supplied by public procurement, the national unified healthcare system (SUS) has promoted access to primary and emergency care and reached universal coverage of vaccination and prenatal care. As this system is increasingly weakened over the decades, the current infrastructure of primary care clinics and emergency units is mainly public. In contrast, hospitals, outpatient clinics, and diagnostic and therapeutic services are mostly private [5, 10]. Consequently, even before a strenuous situation on the health system, such as a pandemic, the diminishing public health expenditures are insufficient and underfunded [5, 4]. Furthermore, the inadequate funds allocated to the system are severely undermined by corruption in public procurement. The diversion of funds meant to attend to social needs contributes to the increase in poverty and inequality [1, 30] in an already unequal country[9]. Even with these setbacks, SUS has been able to attend the population [6, 5, 4], albeit with execution flaws often engendered by the ultimately political challenges the system faces [7, 31].
@@
In 2020, the world faced a pandemic that urged the acquisition of high-demand and essential goods [3, 1], such as personal protection equipment. Consequently, the COVID-19 outbreak highlighted significant inconsistencies in national plans to contain a pandemic and the lack of good governance and corruption prevention [2, 1, 32]. The Brazilian system’s response to the pandemic also faced political challenges [5] such as significant investments in purchasing over 5 million units of an ineffective treatment to the disease, chloroquine [33, 34, 35]. This context only exacerbates a historical problem in public procurement around the globe: an estimated loss of 10% to 25% of global spending is due to corruption [36, 30].
@@
Public procurement in healthcare is a significant risk area for corruption [30], which can occur by bid-rigging, embezzlement of funds, opacity in governance, misuse of power or level position, nepotism and favoritism in the management, petty corruption in levels of service, fraud and theft of medical resources [1]. One such corrupt practice is cartel formation, which submits the government to conspiring agents who violate the objective of biddings, incurring a loss of efficiency in public biddings on price and welfare in varying degrees [37, 30, 1, 2]. Although descriptive analyses can provide information on how susceptible to fraud the Brazilian procurement system is, as shown by the high price variability for some bidding items [38]], we obtain little information on how the companies participating in a bidding process act together to circumvent the inspection system. Also, the collusion between companies in bidding processes is hard to identify and can take multiple forms. For instance, the recurrence of bids for a particular item [2, 30] can contribute to cartel companies’ prior planning and their subsequent organization in rounds and punishment of the players who fail to cooperate in the next bid.
Companies that act in collusion usually participate in a large amount of bidding to share the average earnings or confer an impression of honesty in the whole process [30]. Therefore, each bidding item is a network in which each node is a company participating in the bidding process. Each link exists if two companies are competitors in the same contest. To that end, multiple links represent multiple instances where the companies were co-bidding. As such, the bidding process possesses a complex topology structure that is very informative but often difficult to analyze. More specifically, considering all the bids of a given item, the network is composed of nodes weighted by the number of participating bids, linked by edges weighted by the number of times two companies took part in a bidding process.
@@
We consider data from the Brazilian Federal Government’s public bids from 2013 to 2021 for three personal protective equipment: a surgical mask, rubber glove and endotracheal tubes. The data is freely available through the Transparency Portal [39]. This channel monthly publicizes information gathered from several Federal Government systems. We tested the networks of three products - surgical mask, rubber glove and endotracheal tubes - over six steps. Figure 10 shows the networks for each one of the items, and a more detailed look at the companies year by year can be found at Supplementary Material S1.
@@
Each item’s market has some peculiarities. Masks were in great demand by institutions and individuals during the pandemic period, with a massive increase in demand and offer. Conversely, gloves are items with a relatively predictable annual order, used exclusively by healthcare workers. Despite the pandemic, canceling elective surgeries can correlate to maintaining this market’s stable behavior.
It is important to note that a few companies hugely dominate the public bidding market for these items. Of the 733 companies participating in public bids for endotracheal tubes, 55% won at least one process. However, a surprising 3.5% won half of all submissions. Of the public proposals for masks and gloves, these numbers are 4% and 2%, respectively. Since the internal procedures of Brazilian bidding supervision agencies are - understandably - not publicly available, our approach relies on metrics and methods from covert network analysis, focusing on the sequential removal of edges [24, 25] via the disparity filter.
@@
The most informative subgraph is chosen by maximizing the Heron Coefficient in the network, as illustrated in Figure 11, which depicts the aggregated bidding processes for tubes in 2021. This optimization is performed iteratively, adjusting the significance level $\alpha$ at each step to maximize the network’s Heron Coefficient. Interestingly, the reduction of the significance level implies in an increase of the fraction of winners and, at the sixth step the resulting network usually contains only winners companies.
Figure 12 shows the evolution of the highest significance level of links in the networks, the percentage of winning companies in the network and the common elements in 3 different periods: pre-pandemic (2013-2019), the pandemic’s first year (2020) and its second year (2021). Further details of the significance level’s experiments and results can be checked on the Supplementary Material S1.
@@
The evolution of $\alpha$ with the removal of redundant links is similar for gloves (panels A to C) and endotracheal tubes (panels D to F), both before and after the pandemic. The six steps quickly find a subgraph with relevant links, with $\alpha\leq 10\%$. Concurrently, the proportion of winning companies within these subgraphs increases at each step, indicating that the most significant connections persist among successful bidders. In the tubes network, all companies in the sixth step had won at least one bid. The most significant difference between these two networks emerges in their temporal dynamics. The gloves network exhibits a stable composition, with 2021 closely resembling 2020 and earlier years (2013–2019), reflecting market continuity. Conversely, the tubes network underwent a marked disruption in 2020, followed by a reversion in 2021 toward its pre-2020 structure. Although many firms exited between 2019 and 2020, several long-standing participants reappeared in 2021. Thus, the 2013–2019 data can predict 2021 participants, but not those from the anomalous year of 2020. This evolution illustrates the sensitivity of the Heron metric to shifts in market structure: its earlier peaks captured potential collusive patterns, while the 2020 surge signals renewed concentration that warrants closer monitoring.
@@
In the surgical gloves network, shown in Figure 13, HIC values remain consistently high, but moderate anomalies emerge in 2019 and 2020, marking localized reorganizations rather than systemic disruptions during the pandemic. The smooth upward trend and moderate variability in density and clustering both suggest a gradual structural change, with increasing efficiency achieved through selective, rather than expansive, connectivity. The persistence of multiple components in earlier intervals, followed by consolidation, reinforces the interpretation of slow but steady and selective concentration of the market.
@@
In the endotracheal tubes network, seen in Figure 14, we find a distinct pattern of structural contraction. HIC remains stable at high values through most of the period but declines sharply toward the end, coinciding with a reduction in nodes and edges. The presence of strong positive and negative anomaly scores signals alternating phases of expansion and collapse, suggesting a system characterized by transient structural fragility rather than gradual stabilization. Therefore, even though the general trend is the same as the gloves network, its finer details have significant differences.
@@
Overall, our method captures the significance of the remaining links and elements in the system. Allied with the composition comparison with previous years, this approach composes a detailed characterization of the system and its dynamics. Across all three cases, the inverse relation between HIC and density remains consistent. Moreover, the method’s anomaly detection accurately pinpoints critical transition points, distinguishing the gradual evolution seen with surgical gloves and endotracheal tubes from more abrupt reconfigurations, as seen with masks. These results confirm the robustness and interpretability of our HIC-based temporal framework. Our method effectively captures both smooth and disruptive structural dynamics and identifies critical change intervals, providing a useful tool for covert networks detection.
@@ 5 Conclusion
This work introduces Heron’s Information Coefficient as a practical measure to assess structural imbalances in dynamic networks. By applying it to Brazilian health procurement data spanning the COVID-19 pandemic, we detected patterns that are consistent with collusive behavior and market concentration. The results also show that the method remains useful across synthetic graphs with varying levels of corruption and different configurations.
While promising, this approach should be treated as complementary to existing detection frameworks and not as a stand-alone diagnostic. The method’s geometric intuition helps interpret deviations that conventional measures might smooth over, but it does not capture all institutional or legal nuances. As many studies of covert criminal networks, our work also faces limited access to reliable real-world data for broader validation.
Future studies should compare the Heron coefficient systematically with other dissimilarity or entropy-based metrics and explore its use in near real-time monitoring. We believe this measure can add value to research on procurement integrity, fraud prevention, the resilience of complex systems, as well as, more broadly, community detection.
@@
The way the vertices interact and connect provides essential information about the system. In particular, information about their connectivity patterns represents a valuable tool on how different dynamics occur in the system. For example, the speed at which information travels within a cohesive community is much greater than in a more sparse one. The ideal public bidding network consists of a network where nodes are connected uniformly. For example, in a total of 100 biddings, if a company $x$ participates of 10 ($n_{x}=10$) and a company $y$ participates of 15 ($n_{y}=15$), it is expected that the probability that these two companies participate in the same bidding is $0.1\times 0.15$, resulting in an average of 1.5 joint participations. Values high above 1.5 must indicate some abnormal relationships. Here, we identify the evolution of links happening in a randomly uniform (Poisson) manner and links that occur with a higher probability than the expected [40].
@@
To capture how different a random network composed of Poisson-like elements is from the real interaction between the elements, we built a $G_{P}$ network (Poisson network) from the fact that the probability of a link existing between two companies depends only on the number of times they participate in a contest. This network is compared with the real $G_{R}$ network where the probability of a link occurring is given by the fraction of its weight ($w$) by the total number of bids. More specifically, be $G=(V,E,W_{v},W_{e})$ the bidding network of an item where $V$ is the set of vertices, $E$ the set of links, $W_{v}$ the set of the number of bids that each vertex participated and $W_{e}$ the number of times the link $e$ occurred, $G_{P}=(V,E)$ the unweighted, where the probability of a link occurring is Poisson and $G_{R}$ the real network. We then compare the Heron’s values for these two different networks generations. Table S1 compares the real bidding network with a Poisson one, showcasing how our coefficient could be used in a yearly comparison. In practical terms, the competent authorities should further investigate a deviating behavior because the real-world network’s dynamics in bidding processes do not possess a uniformly random distribution.

### v2

@@ 4 Experimental Results
All the results on this section are based on the extraction of the network’s backbone with a disparity filter based on HIC maximization. The source code for the characterization of our coefficient’s behavior and other experiments can be found on GitHub (https://github.com/FutureLab-DCC/Heron_coefficient).
@@ 4.1 Artificial bidding network
We tested our method in artificial procurement networks generated using the Bernoulli Weighted Random Network (BWRN) model [14], which combines the Stochastic Block Model (SBM) for group formation with hierarchical structures for management layers. This generator creates synthetic networks that are statistically equivalent substitutes for real networks, allowing for research on covert networks. Its validity has been established through comparison with the publicly available Sicilian Mafia network [9], which is a covert network that covers phone calls and in-person meetings extracted from court proceedings. Thus, these synthetic networks replicate the behaviors characteristic of real colluding agents while addressing a critical challenge in covert network research: data availability [8]. We enhanced realism by incorporating successful bids to better simulate procurement networks with known corrupt agents. Then, we iterated the HIC filtering process until network sparsity prevented further analysis.
@@
To validate robustness, we tested the method in 100 random seeds and varying parameters, iterating until all HIC values reached zero, that is, the network dissolved. The results were consistent, as seen in 9. With a lower percentage of corrupt agents, the method stopped faster. Networks with lower corruption levels (5%, 10%) maintained higher accuracy throughout, generally staying above 0.8. Furthermore, the method showed great final accuracy for networks containing up to 30% corrupt agents, with the final iteration successfully identifying at least one corrupt agent in most cases. Performance remains stable across network scales, although optimization would be necessary for graphs exceeding 10,000 nodes.
@@
In summary, our method is computationally cheap, converges fast, and shows promising results in detecting corrupt agents in bidding processes. The Heron Information Coefficient effectively identifies corrupted agents in early iterations, particularly at lower corruption levels. While accuracy becomes less stable at corruption levels exceeding 50% - an expected limitation since anomaly detection inherently struggles when anomalous nodes constitute the majority - the method remains valuable. Even after six iterations in high-corruption scenarios, it substantially reduces network size while retaining corrupt nodes. As such, our results suggest that the Heron coefficient can serve as a practical aid for corruption inspection, even in resource-constrained environments with limited detection capabilities.
@@ 4.2 The BID Network
Brazil is a country of continental dimensions, the fifth largest in the world and the sixth in terms of population [21] with considerable regional and social inequalities [13, 5, 30]. As determined by the country’s Constitution, the provision of services in healthcare and education is a citizen’s right, met with resources collected through taxation at the federal, state, and municipal levels [30, 18, 13, 39]. Supplied by public procurement, the national unified healthcare system (SUS) has promoted access to primary and emergency care and reached universal coverage of vaccination and prenatal care. As this system has become increasingly weak over the decades, the current infrastructure of primary care clinics and emergency units is mainly public. In contrast, hospitals, outpatient clinics, and diagnostic and therapeutic services are mostly private [31, 13]. Consequently, even before a strenuous situation on the health system, such as a pandemic, the decreasing public health expenditures are insufficient and underfunded [31, 30]. Furthermore, the inadequate funds allocated to the system are severely undermined by corruption in public procurement. The diversion of funds meant to address social needs contributes to the increase in poverty and inequality [37, 38] in an already unequal country[5]. Even with these setbacks, SUS has been able to attend to the population [18, 31, 30], albeit with execution flaws often engendered by the ultimately political challenges the system faces [39, 27].
@@
In 2020, the world faced a pandemic that urged the acquisition of essential and high-demand goods [28, 37], such as personal protection equipment. Consequently, the COVID-19 outbreak highlighted significant inconsistencies in national plans to contain a pandemic and the lack of good governance and corruption prevention [35, 37, 22]. The response of the Brazilian system to the pandemic also faced political challenges [31] such as significant investments in the purchase of more than 5 million units of an ineffective treatment of the disease, chloroquine [42, 1, 12]. This context only exacerbates a historical problem in public procurement around the world: an estimated loss of 10% to 25% of global spending is due to corruption [24, 38].
@@
Public procurement in healthcare is a significant area of risk of corruption [38], which can occur through bid-rigging, funds embezzlement, opacity in governance, misuse of power, nepotism and favoritism in management, petty corruption in levels of service, fraud and theft of medical resources [37]. One such corrupt practice is cartel formation, which exposes the government to conspiring agents who violate the objective of bidding, incurring a loss of efficiency in public biddings on price and welfare in varying degrees [20, 38, 37, 35]. Although descriptive analyzes can provide information on how susceptible to fraud the Brazilian procurement system is, as shown by the high price variability for some bidding items [19]], we obtain little information on how the companies participating in a bidding process act together to bypass the inspection system.
@@
We consider data from the Brazilian Federal Government’s public bids from 2013 to 2021 for three personal protective equipment: a surgical mask, rubber gloves and endotracheal tubes. The data is freely available through the Transparency Portal [10]. This channel monthly publicizes information gathered from several Federal Government systems. Considering all the bids of a given item, the network is composed of nodes weighted by the number of participating bids, linked by edges weighted by the number of times two companies took part in a bidding process. We tested the networks of these three products over six steps. Figure 12 shows the networks for each one of the items, and a more detailed look at the companies year by year can be found in the Supplementary Material S2.
@@
The market for each item has some peculiarities. Masks were in great demand by institutions and individuals during the pandemic period, with a massive increase in demand and offer. Conversely, gloves are items with a relatively predictable annual order, used exclusively by healthcare workers. Despite the pandemic, canceling elective surgeries can correlate to maintaining this market’s stable behavior.
It is important to note that a few companies hugely dominate the public bidding market for these items. Of the 733 companies participating in public bids for endotracheal tubes, 55% won at least one process. However, a surprising 3.5% won half of all submissions. Of the public proposals for masks and gloves, these numbers are 4% and 2%, respectively. Since the internal procedures of Brazilian bidding supervision agencies are, understandably, not publicly available, our approach relies on metrics and methods from covert network analysis, focusing on the sequential removal of edges [9, 16] via the disparity filter.
@@
The most informative subgraph is chosen by maximizing the Heron Coefficient in the network, as illustrated in Figure 13, which depicts the aggregate bidding processes for tubes in 2021. This optimization is performed iteratively, adjusting the significance level $\alpha$ at each step to maximize the network’s Heron Coefficient. Interestingly, the reduction of the significance level implies in an increase of the fraction of winners and, at the sixth step the resulting network usually contains only winners companies.
Figure 14 shows the evolution of the highest significance level of links in the networks, the percentage of winning companies in the network and the common elements in 3 different periods: pre-pandemic (2013-2019), the pandemic’s first year (2020) and its second year (2021). Further details of the significance level’s experiments and results can be checked on the Supplementary Material S2.
@@
The evolution of $\alpha$ with the removal of redundant links is similar for gloves (panels A to C) and endotracheal tubes (panels D to F), both before and after the pandemic. The six steps quickly find a subgraph with relevant links, with $\alpha\leq 10\%$. Concurrently, the proportion of winning companies within these subgraphs increases at each step, indicating that the most significant connections persist among successful bidders. In the tubes network, all companies in the sixth step had won at least one bid. The most significant difference between these two networks emerges in their temporal dynamics. The gloves network exhibits a stable composition, with 2021 closely resembling 2020 and earlier years (2013–2019), reflecting market continuity. In contrast, the tubes network underwent a marked disruption in 2020, followed by a reversion in 2021 toward its pre-2020 structure. Although many firms exited the market between 2019 and 2020, several long-standing participants reappeared in 2021. Thus, the data for 2013–2019 can predict 2021 participants, but not those from the anomalous year of 2020. This evolution illustrates the sensitivity of the Heron metric to shifts in market structure: its earlier peaks captured potential collusive patterns, while the 2020 surge signals renewed concentration that warrants closer monitoring.
@@
In the surgical gloves network, shown in Figure 15, HIC values remain consistently high, but moderate anomalies emerge in 2019 and 2020, marking localized reorganizations rather than systemic disruptions during the pandemic. The smooth upward trend and moderate variability in density and clustering both suggest a gradual structural change, with increasing efficiency achieved through selective, rather than expansive, connectivity. The persistence of multiple components in earlier intervals, followed by consolidation, reinforces the interpretation of slow but steady and selective market concentration.
@@
In the network of endotracheal tubes, seen in Figure 16, we find a distinct pattern of structural contraction. HIC remains stable at high values for most of the period, but decreases sharply toward the end, coinciding with a reduction in nodes and edges. The presence of strong positive and negative anomaly scores signals alternating phases of expansion and collapse, suggesting a system characterized by transient structural fragility rather than gradual stabilization. Therefore, even though the general trend is the same as the gloves network, its finer details have significant differences.
@@
Overall, our method captures the significance of the remaining links and elements in the system. Allied with the composition comparison with previous years, this approach composes a detailed characterization of the system and its dynamics. Across all three cases, the inverse relationship between HIC and density remains consistent. Moreover, the method’s anomaly detection accurately pinpoints critical transition points, distinguishing the gradual evolution seen with surgical gloves and endotracheal tubes from more abrupt reconfigurations, as seen with masks. These results confirm the robustness and interpretability of our HIC-based temporal framework. Our method effectively captures both smooth and disruptive structural dynamics and identifies critical change intervals, providing a useful tool for covert networks detection.
This work establishes Heron’s Information Coefficient (HIC) as a robust metric for quantifying structural imbalances within dynamic networks. Applied to Brazilian health procurement data during the COVID-19 pandemic, HIC successfully isolated topological anomalies consistent with collusive behavior and market concentration—findings further corroborated by stress-testing on synthetic graphs with varying corruption densities. While the method’s geometric intuition effectively exposes deviations often smoothed over by conventional statistics, we emphasize that it functions best as a complementary diagnostic rather than a standalone proof of illicit activity, particularly given the inherent scarcity of ground-truth criminal data. Future research should focus on benchmarking HIC against established information-theoretic metrics and developing real-time monitoring pipelines, thereby broadening its applicability to procurement integrity, fraud detection, and the resilience of complex systems.
@@ 5 Conclusion
This work introduces Heron’s Information Coefficient as a robust measure to quantify structural imbalances in dynamic networks. Applied to Brazilian health procurement data spanning the COVID-19 pandemic, HIC successfully isolated topological anomalies consistent with collusive behavior and market concentration - findings further corroborated by stress-testing on synthetic graphs with varying corruption densities and an aggregate procurement dataset.
While promising, this approach should be treated as complementary to existing detection frameworks and not as a stand-alone diagnostic. The method’s geometric intuition helps interpret deviations that conventional measures might smooth over, but it does not capture all institutional or legal nuances. As in many studies of covert criminal networks, our work also faces limited access to reliable real-world data for broader validation. We emphasize that our method functions best as a complementary diagnostic rather than a standalone proof of illicit activity, particularly given the inherent scarcity of ground-truth criminal data and the ever-evolving dynamics of criminal activity.
Future studies should systematically compare the Heron coefficient with other dissimilarity or entropy-based metrics and explore its use in other anomaly detection scenarios. We believe that this measure can add value to research on procurement integrity, fraud prevention, the resilience of complex systems, as well as, more broadly, community detection.
Appendix S1 Complexity analysis
In its standard formulation, the algorithm exhibits a time complexity of $\mathcal{O}(|V|(|V|+|E|))$ and a space complexity of $\mathcal{O}(|V|^{2})$. While these bounds are manageable for small-to-medium networks, they present significant scaling challenges for graphs exceeding $10^{4}$ nodes, primarily due to memory constraints. The performance cost is driven by the Network Node Dispersion (NND) function. Computing the topological dispersion requires an All-Pairs Shortest Path (APSP) calculation. For unweighted graphs, this is implemented via Breadth-First Search (BFS) traversals from each node, yielding a time complexity of $\mathcal{O}(|V|(|V|+|E|))$. More critically, the naive storage of the resulting distance matrix necessitates $\mathcal{O}(|V|^{2})$ space, creating a dense memory bottleneck even for sparse input graphs.
This cost is compounded by the iterative Heron optimization loop. To identify the optimal cut, the algorithm sweeps through $K$ quantile steps (typically $K=100$), requiring dual NND evaluations at each step. If the pruning process continues for $M$ generations, the total theoretical complexity scales to $\mathcal{O}(M\cdot K\cdot|V|\cdot|E|)$.
A critical design choice in our framework is the exclusion of the alpha-centrality term ($w_{3}=0$) for large-scale analysis. Including this term requires operations on the graph complement - transforming sparse inputs into dense representations - which would elevate the time complexity to $\mathcal{O}(|V|^{3})$. While the third term enhances sensitivity to spectral isomorphism, its exclusion is a necessary and acceptable approximation for backbone extraction. In this context, preserving the dominant connectivity patterns (captured by the first two terms) takes precedence over the detection of subtle spectral symmetries, rendering the cubic computational cost unjustifiable.
To overcome the $\mathcal{O}(|V|^{2})$ memory barrier, we implemented a batched parallelization strategy. Because the graph topology evolves sequentially during pruning, caching is ineffective. Instead, we utilize:
  1. 1.
    Batched Pathfinding: Rather than allocating a monolithic distance matrix, we compute shortest paths in strip-wise batches. Statistics (entropy and histograms) are aggregated on-the-fly, and raw distance data is immediately discarded. This reduces memory usage to $\mathcal{O}(|V|\times\text{batch\_size})$.
  2. 2.
    Integrated Diameter Calculation: The network diameter is derived dynamically during the pathfinding batch, eliminating the need for a separate $\mathcal{O}(|V||E|)$ traversal.
  3. 3.
    Parallel Execution: The batching process allows for efficient distribution across multi-core architectures, providing near-linear speedup in the NND calculation phase.
This optimized approach shifts the bottleneck from memory capacity to CPU throughput, making the method viable for larger systems without sacrificing the rigorous topological analysis required by the disparity filter.
@@
The way the vertices interact and connect provides essential information about the system. In particular, information about their connectivity patterns represents a valuable tool on how different dynamics occur in the system. For example, the speed at which information travels within a cohesive community is much greater than in a more sparse one. The theoretical benchmark of public bidding network consists of a network where nodes are connected uniformly. For example, in a total of 100 biddings, if a company $x$ participates in 10 ($n_{x}=10$) and a company $y$ participates of 15 ($n_{y}=15$), it is expected that the probability that these two companies participate in the same bidding is $0.1\times 0.15$, resulting in an average of 1.5 joint participation. Values above 1.5 must indicate some abnormal relationships. Here, we identify the evolution of links happening in a random uniform manner and links that occur with a higher probability than the expected null hypothesis [2].
@@
To capture how different a random network composed of Poisson-like elements is from the real interaction between the elements, we built a $G_{P}$ network (Poisson network) from the fact that the probability of a link existing between two companies depends only on the number of times they participate in a contest. This network is compared with the real $G_{R}$ network where the probability of a link occurring is given by the fraction of its weight ($w$) by the total number of bids. More specifically, let $G=(V,E,W_{v},W_{e})$ be the bidding network of an item where $V$ is the set of vertices, $E$ the set of links, $W_{v}$ the set of the number of bids that each vertex participated in and $W_{e}$ the number of times the link $e$ occurred, $G_{P}=(V,E)$ the unweighted, where the probability of a link occurring is Poisson and $G_{R}$ the real network. We then compare the Heron’s values for these two different networks generations. Table S1 presents the evolution of $H_{real}$ versus the expected $H_{null}$. We report the ratio $H_{real}/H_{null}$ to illustrate the magnitude of structural deviation. Bold values indicate statistical significance ($p<0.05$), determined via a Z-test against the null ensemble distribution. A ratio significantly greater than 1.0 implies that the information geometry of the real network is fundamentally more complex than what can be explained by simple participation rates, warranting further investigation by competent authorities.